OpenStudy (anonymous):

Maria throws an apple vertically upward from a height of 1.3 m with an initial velocity of + 2.4 m/s. a. Will the apple reach a friend in a treehouse 5.3 m above the ground? b. If the apple is not caught, how long will the apple be in the air before it hits the ground?

OpenStudy (anonymous):

Find the maximum height apple will reach first: \[v^2=v_0^2-2gh \\v=0 \text{ since it stopping at highest point} \\0=v_0^2-2gh\\h=\frac{v_0^2}{2g} \] Can you find h?